# An Algebraic Roadmap of Particle Theories Part I: General construction

### v1

@@ IV.6 Division algebraic reflections
It is straightforward to identify analog reflections in the subsequent Clifford algebraic factors of equation (24). For example, a $\mathbb{Z}_{2}$-grading on $\mathcal{L}_{\mathbb{H}}\simeq Cl(0,2)$ may be induced by a quaternionic reflection *$*_{\overrightarrow{\mathbb{H}}}:L_{\epsilon_{m}}\mapsto-L_{\epsilon_{m}}$ for $m\in\{1,2\}$. Similarly, a $\mathbb{Z}_{2}$-grading on $\mathcal{L}_{\mathbb{O}}\simeq Cl(0,6)$ may be induced by an octonionic reflection *$*_{\overrightarrow{\mathbb{O}}}:L_{e_{j}}\mapsto-L_{e_{j}}$ for $j\in\{1,2,\dots 6\}$.
@@ IV.7 An explicit division algebraic representation of $\mathbb{C}l(10)$
$\begin{array}[]{lll}a_{1}^{\dagger}:=\frac{1}{2}(\Gamma_{5}+i\Gamma_{4})&&a_{1}:=\frac{1}{2}(-\Gamma_{5}+i\Gamma_{4})\\ a_{2}^{\dagger}:=\frac{1}{2}(\Gamma_{3}+i\Gamma_{1})&&a_{2}:=\frac{1}{2}(-\Gamma_{3}+i\Gamma_{1})\\ a_{3}^{\dagger}:=\frac{1}{2}(\Gamma_{6}+i\Gamma_{2})&&a_{3}:=\frac{1}{2}(-\Gamma_{6}+i\Gamma_{2})\\ a_{4}^{\dagger}:=\frac{1}{2}(\Gamma_{8}+i\Gamma_{7})&&a_{4}:=\frac{1}{2}(-\Gamma_{8}+i\Gamma_{7})\\ a_{5}^{\dagger}:=\frac{1}{2}(\Gamma_{10}+i\Gamma_{9})&&a_{5}:=\frac{1}{2}(-\Gamma_{10}+i\Gamma_{9}),\\ \end{array}$ (28)
@@ IV.8 Cascade of particle symmetries:
$\begin{array}[]{ll}\ell_{10}^{D}:=&r_{ij}L_{e_{i}}L_{e_{j}}+r_{m}L_{\epsilon_{m}}+r^{\prime}_{m}L_{\epsilon_{m}}L_{e_{7}}R_{\epsilon_{3}}\\ &+\hskip 2.84526ptr_{mj}^{\prime}iL_{\epsilon_{m}}L_{e_{j}}+r_{j}^{\prime\prime}iL_{e_{j}}L_{e_{7}}R_{\epsilon_{3}}\end{array}$ (32)
@@
$\begin{array}[]{lc}\ell_{10}^{D}\Psi_{\textup{L}}=&(r_{ab}L_{e_{a}}L_{e_{b}}+r^{\prime}_{mn}L_{\epsilon_{m}}L_{\epsilon_{n}}+r_{ma}^{\prime\prime}iL_{\epsilon_{m}}L_{e_{a}})\Psi_{\textup{L}},\\ &\mathfrak{spin}(10)\end{array}$ (33)
@@
$\begin{array}[]{lc}\ell_{10}:=&r_{ab}L_{e_{a}}L_{e_{b}}+r^{\prime}_{mn}L_{\epsilon_{m}}L_{\epsilon_{n}}+r_{ma}^{\prime\prime}iL_{\epsilon_{m}}L_{e_{a}}.\\ &\mathfrak{spin}(10)\end{array}$ (34)
@@
$\begin{array}[]{rll}{\ell}_{\textup{PS}}\Psi_{\textup{L}}=(r_{ij}L_{e_{i}}L_{e_{j}}&+&r_{m}L_{\epsilon_{m}}+r^{\prime}_{m}iL_{\epsilon_{m}}L_{e_{7}})\Psi_{\textup{L}}\\ \mathfrak{spin}(6)&\oplus&\mathfrak{spin}(4)\end{array}$ (36)
@@
Next, we will consider those symmetries furthermore invariant under an analogous octonionic reflection, deriving from right multiplication. Namely *$*_{\overleftarrow{\mathbb{O}}}:R_{e_{j}}\mapsto-R_{e_{j}}$ $\forall j\in\{1,\dots,6\}.$
@@
$\begin{array}[]{rll}{\ell}_{\textup{LR}}\Psi_{\textup{L}}=(r_{ij}^{\prime\prime}L_{e_{i}}L_{e_{j}}&+&r_{m}L_{\epsilon_{m}}+r^{\prime}_{m}iL_{\epsilon_{m}}L_{e_{7}})\Psi_{\textup{L}},\\ \mathfrak{u}(3)&\oplus&\mathfrak{spin}(4)\end{array}$ (38)
@@
$\begin{array}[]{lll}i\Lambda_{1}:=\frac{1}{2}\left(L_{34}-L_{15}\right)&&i\Lambda_{2}:=\frac{1}{2}\left(L_{14}+L_{35}\right)\\ i\Lambda_{3}:=\frac{1}{2}\left(L_{13}-L_{45}\right)&&i\Lambda_{4}:=-\frac{1}{2}\left(L_{25}+L_{46}\right)\\ i\Lambda_{5}:=\frac{1}{2}\left(L_{24}-L_{56}\right)&&i\Lambda_{6}:=-\frac{1}{2}\left(L_{16}+L_{23}\right)\\ i\Lambda_{7}:=-\frac{1}{2}\left(L_{12}+L_{36}\right)&&i\Lambda_{8}:=\frac{-1}{2\sqrt{3}}\left(L_{13}+L_{45}-2L_{26}\right),\end{array}$ (39)
@@
An alternative route was found by Boyle in [28] that maps Spin(10) directly to the Left-Right Symmetric model via *$*_{\overleftarrow{\mathbb{O}}}$ in the context of the exceptional Jordan algebra. This result likewise holds in our current model.
@@
$\begin{array}[]{rllrll}{\ell}_{{\textup{UBZ+}}}\Psi_{\textup{L}}=(r_{ij}^{\prime\prime}L_{e_{i}}L_{e_{j}}&+&r_{3}L_{\epsilon_{3}}&+&r^{\prime}_{3}iL_{\epsilon_{3}}L_{e_{7}})\Psi_{\textup{L}}.\\ \mathfrak{u}(3)&\oplus&\mathfrak{u}(1)&\oplus&\mathfrak{u}(1)\end{array}$ (42)
@@
$\begin{array}[]{rll}{\ell}_{{\textup{UB+}}}\Psi_{\textup{L}}=(r_{ij}^{\prime\prime}L_{e_{i}}L_{e_{j}}&+&r_{3}L_{\epsilon_{3}})\Psi_{\textup{L}},\\ \mathfrak{u}(3)&\oplus&\mathfrak{u}(1)\end{array}$ (44)
@@ IV.9 Cascade of particle symmetries: the Georgi-Glashow line
$\begin{array}[]{lll}\ell_{5}&=&R_{cd}\left(a_{c}a^{\dagger}_{d}-a_{d}a^{\dagger}_{c}\right)\\ &+&R_{cd}^{\prime}\hskip 1.42262pti\left(a_{c}a^{\dagger}_{d}+a_{d}a^{\dagger}_{c}\right)\\ &+&R_{c}\hskip 1.42262pti\left(a_{c}a^{\dagger}_{c}-a_{c+1}a^{\dagger}_{c+1}\right),\end{array}$ (45)
@@
$\begin{array}[]{lllll}&&\ell_{5}\Psi_{\textup{L}}=&&\\ &&\big{(}r_{j}i\Lambda_{j}+r^{\prime}_{m}L_{\epsilon_{m}}s+irY&&\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{1}\left(L_{\epsilon_{1}}L_{e_{4}}+L_{\epsilon_{2}}L_{e_{5}}\right)&+&it_{1}^{\prime}\left(L_{\epsilon_{1}}L_{e_{5}}-L_{\epsilon_{2}}L_{e_{4}}\right)\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{1}^{\prime\prime}\left(L_{\epsilon_{3}}L_{e_{4}}-iL_{e_{5}}L_{e_{7}}\right)&+&it_{1}^{\prime\prime\prime}(L_{\epsilon_{3}}L_{e_{5}}+iL_{e_{4}}L_{e_{7}})\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{2}\left(L_{\epsilon_{1}}L_{e_{1}}+L_{\epsilon_{2}}L_{e_{3}}\right)&+&it_{2}^{\prime}(L_{\epsilon_{1}}L_{e_{3}}-L_{\epsilon_{2}}L_{e_{1}})\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{2}^{\prime\prime}\left(L_{\epsilon_{3}}L_{e_{1}}-iL_{e_{3}}L_{e_{7}}\right)&+&it_{2}^{\prime\prime\prime}(L_{\epsilon_{3}}L_{e_{3}}+iL_{e_{1}}L_{e_{7}})\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{3}\left(L_{\epsilon_{1}}L_{e_{2}}+L_{\epsilon_{2}}L_{e_{6}}\right)&+&it_{3}^{\prime}(L_{\epsilon_{1}}L_{e_{6}}-L_{\epsilon_{2}}L_{e_{2}})\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{3}^{\prime\prime}\left(L_{\epsilon_{3}}L_{e_{2}}-iL_{e_{6}}L_{e_{7}}\right)&+&it_{3}^{\prime\prime\prime}(L_{\epsilon_{3}}L_{e_{6}}+iL_{e_{2}}L_{e_{7}})\big{)}\Psi_{\textup{L}}\\ &&\hskip 48.36958pt\mathfrak{su}(5)&&\end{array}$ (46)
@@
With our $\mathfrak{su}(5)$ action now established, division algebraic reflections may be deployed so as to break these symmetries sequentially.
@@
$\begin{array}[]{rccll}\ell_{\textup{SM}}\Psi_{\textup{L}}=\big{(}r_{j}i\Lambda_{j}&+&r^{\prime}_{m}L_{\epsilon_{m}}s&+&irY\big{)}\Psi_{\textup{L}}.\\ \mathfrak{su}(3)&\oplus&\mathfrak{su}(2)&\oplus&\mathfrak{u}(1)\end{array}$ (50)
@@
$\begin{array}[]{rccll}\ell_{\textup{UBZ}}\Psi_{\textup{L}}=\big{(}r_{j}i\Lambda_{j}&+&r^{\prime}_{3}L_{\epsilon_{3}}s&+&irY\big{)}\Psi_{\textup{L}}.\\ \mathfrak{su}(3)&\oplus&\mathfrak{u}(1)&\oplus&\mathfrak{u}(1)\end{array}$ (52)
@@
$\begin{array}[]{lrcl}\ell_{\textup{UB}}\Psi_{\textup{L}}=&\big{(}r_{j}i\Lambda_{j}&+&r^{\prime}iQ\big{)}\Psi_{\textup{L}},\\ &\mathfrak{su}(3)&\oplus&\mathfrak{u}(1)\end{array}$ (54)
@@
Particularly satisfying about this last result is the fact that the familiar complex conjugate is what dictates the Standard Model’s final unbroken gauge symmetries. Conceivably, the Standard Model’s gauge group could have broken in a different way, had it been paired with a different Higgs sector. Could the complex conjugate ultimately be stewarding the symmetry breaking process?
@@ V Roadmap
In this paper, we have listed five different forms of symmetry breaking steps. Namely, the multivector condition, $\Psi_{V}$, invariance under two types of octonionic reflection, *${*_{\overrightarrow{\mathbb{O}}}}$ and *$*_{\overleftarrow{\mathbb{O}}}$, one quaternionic reflection, $*_{\overrightarrow{\mathbb{H}}},\vspace{.5mm}$ and one complex reflection, $*_{\overrightarrow{\mathbb{C}}}$. These four types of reflection are precisely those that can be realized in terms of the left-multiplication algebra of $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}.$
@@
Our culminating diagram results when all five symmetry breaking steps are amalgamated ($\Psi_{V}$, *${*_{\overrightarrow{\mathbb{O}}}}$, *$*_{\overleftarrow{\mathbb{O}}}$, $*_{\overrightarrow{\mathbb{H}}},$$*_{\overrightarrow{\mathbb{C}}}$). For reflections, we begin with the octonions, continue to the quaternions, and close with the complex numbers. Please see Figure (6) at the end of this article.
Figure 6: Six well-known models of elementary particles are interconnected algebraically into a detailed particle roadmap. The “SO(10)” model (Spin(10)), the Georgi-Glashow model (SU(5)), the Pati-Salam model (PS), the Left-Right Symmetric model (LR), the Standard Model pre-Higgs-mechanism (SM), and the Standard Model post-Higgs-mechanism (UB) are each interrelated via certain algebraic constraints. The multivector condition is represented by edges labeled as $\Psi_{V}$. Division algebraic reflection constraints are represented by edges labeled as *$*_{\mathbb{D}}$. Beyond these six well-known models, there are three more. For the definition of UBZ+ see equation (42); for the definition of UB+ see equation (44); for the definition of UBZ see equation (52).
From the outset, readers may notice an interesting feature of this 5-step network. Namely, not all pathways are unique; a number of them bifurcate.
As a first example, we identify two compatible symmetry breaking pathways from the Spin(10) model to the Left-Right Symmetric model. One may first restrict Spin(10) to those symmetries invariant under the octonionic reflection *$*_{\overrightarrow{\mathbb{O}}}$, resulting in the Pati-Salam model. Subsequently imposing invariance under *${*_{\overleftarrow{\mathbb{O}}}}$ then results in the Left-Right Symmetric model. As an alternate route, readers can bypass directly to the Left-Right Symmetric model by simply restricting Spin(10) once via *$*_{\overleftarrow{\mathbb{O}}}$, as in [28].
@@
VI Conclusion

### v2

@@ IV.6 Division algebraic reflections
It is straightforward to identify analog reflections in the subsequent Clifford algebraic factors of equation (24). For example, a $\mathbb{Z}_{2}$-grading on $\mathcal{L}_{\mathbb{H}}\simeq Cl(0,2)$ may be induced by a quaternionic reflection $*_{\overrightarrow{\mathbb{H}}}:L_{\epsilon_{m}}\mapsto-L_{\epsilon_{m}}$ for $m\in\{1,2\}$. Similarly, a $\mathbb{Z}_{2}$-grading on $\mathcal{L}_{\mathbb{O}}\simeq Cl(0,6)$ may be induced by an octonionic reflection $*_{\overrightarrow{\mathbb{O}}}:L_{e_{j}}\mapsto-L_{e_{j}}$ for $j\in\{1,2,\dots 6\}$.
@@ IV.7 An explicit division algebraic representation of $\mathbb{C}l(10)$
$\begin{array}[]{lll}a_{1}^{\dagger}:=\frac{1}{2}(\Gamma_{5}+i\Gamma_{4})&&a_{1}:=\frac{1}{2}(-\Gamma_{5}+i\Gamma_{4})\vspace{2mm}\\ a_{2}^{\dagger}:=\frac{1}{2}(\Gamma_{3}+i\Gamma_{1})&&a_{2}:=\frac{1}{2}(-\Gamma_{3}+i\Gamma_{1})\vspace{2mm}\\ a_{3}^{\dagger}:=\frac{1}{2}(\Gamma_{6}+i\Gamma_{2})&&a_{3}:=\frac{1}{2}(-\Gamma_{6}+i\Gamma_{2})\vspace{2mm}\\ a_{4}^{\dagger}:=\frac{1}{2}(\Gamma_{8}+i\Gamma_{7})&&a_{4}:=\frac{1}{2}(-\Gamma_{8}+i\Gamma_{7})\vspace{2mm}\\ a_{5}^{\dagger}:=\frac{1}{2}(\Gamma_{10}+i\Gamma_{9})&&a_{5}:=\frac{1}{2}(-\Gamma_{10}+i\Gamma_{9}),\vspace{2mm}\\ \end{array}$ (28)
@@ IV.8 Cascade of particle symmetries:
$\begin{array}[]{ll}\ell_{10}^{D}:=&r_{ij}L_{e_{i}}L_{e_{j}}+r_{m}L_{\epsilon_{m}}+r^{\prime}_{m}L_{\epsilon_{m}}L_{e_{7}}R_{\epsilon_{3}}\vspace{2mm}\\ &+\hskip 2.84526ptr_{mj}^{\prime}iL_{\epsilon_{m}}L_{e_{j}}+r_{j}^{\prime\prime}iL_{e_{j}}L_{e_{7}}R_{\epsilon_{3}}\end{array}$ (32)
@@
$\begin{array}[]{lc}\ell_{10}^{D}\Psi_{\textup{L}}=&(r_{ab}L_{e_{a}}L_{e_{b}}+r^{\prime}_{mn}L_{\epsilon_{m}}L_{\epsilon_{n}}+r_{ma}^{\prime\prime}iL_{\epsilon_{m}}L_{e_{a}})\Psi_{\textup{L}},\vspace{2mm}\\ &\mathfrak{spin}(10)\end{array}$ (33)
@@
$\begin{array}[]{lc}\ell_{10}:=&r_{ab}L_{e_{a}}L_{e_{b}}+r^{\prime}_{mn}L_{\epsilon_{m}}L_{\epsilon_{n}}+r_{ma}^{\prime\prime}iL_{\epsilon_{m}}L_{e_{a}}.\vspace{2mm}\\ &\mathfrak{spin}(10)\end{array}$ (34)
@@
$\begin{array}[]{rll}{\ell}_{\textup{PS}}\Psi_{\textup{L}}=(r_{ij}L_{e_{i}}L_{e_{j}}&+&r_{m}L_{\epsilon_{m}}+r^{\prime}_{m}iL_{\epsilon_{m}}L_{e_{7}})\Psi_{\textup{L}}\vspace{2mm}\\ \mathfrak{spin}(6)&\oplus&\mathfrak{spin}(4)\end{array}$ (36)
@@
Next, we will consider those symmetries furthermore invariant under an analogous octonionic reflection, deriving from right multiplication. Namely $*_{\overleftarrow{\mathbb{O}}}:R_{e_{j}}\mapsto-R_{e_{j}}$ $\forall j\in\{1,\dots,6\}.$
@@
$\begin{array}[]{rll}{\ell}_{\textup{LR}}\Psi_{\textup{L}}=(r_{ij}^{\prime\prime}L_{e_{i}}L_{e_{j}}&+&r_{m}L_{\epsilon_{m}}+r^{\prime}_{m}iL_{\epsilon_{m}}L_{e_{7}})\Psi_{\textup{L}},\vspace{2mm}\\ \mathfrak{u}(3)&\oplus&\mathfrak{spin}(4)\end{array}$ (38)
@@
$\begin{array}[]{lll}i\Lambda_{1}:=\frac{1}{2}\left(L_{34}-L_{15}\right)&&i\Lambda_{2}:=\frac{1}{2}\left(L_{14}+L_{35}\right)\vspace{2mm}\\ i\Lambda_{3}:=\frac{1}{2}\left(L_{13}-L_{45}\right)&&i\Lambda_{4}:=-\frac{1}{2}\left(L_{25}+L_{46}\right)\vspace{2mm}\\ i\Lambda_{5}:=\frac{1}{2}\left(L_{24}-L_{56}\right)&&i\Lambda_{6}:=-\frac{1}{2}\left(L_{16}+L_{23}\right)\vspace{2mm}\\ i\Lambda_{7}:=-\frac{1}{2}\left(L_{12}+L_{36}\right)&&i\Lambda_{8}:=\frac{-1}{2\sqrt{3}}\left(L_{13}+L_{45}-2L_{26}\right),\end{array}$ (39)
@@
An alternative route was found by Boyle in [32] that maps Spin(10) directly to the Left-Right Symmetric model via $*_{\overleftarrow{\mathbb{O}}}$ in the context of the exceptional Jordan algebra. This result likewise holds in our current model.
@@
$\begin{array}[]{rllrll}{\ell}_{{\textup{UBZ+}}}\Psi_{\textup{L}}=(r_{ij}^{\prime\prime}L_{e_{i}}L_{e_{j}}&+&r_{3}L_{\epsilon_{3}}&+&r^{\prime}_{3}iL_{\epsilon_{3}}L_{e_{7}})\Psi_{\textup{L}}.\vspace{2mm}\\ \mathfrak{u}(3)&\oplus&\mathfrak{u}(1)&\oplus&\mathfrak{u}(1)\end{array}$ (42)
@@
$\begin{array}[]{rll}{\ell}_{{\textup{UB+}}}\Psi_{\textup{L}}=(r_{ij}^{\prime\prime}L_{e_{i}}L_{e_{j}}&+&r_{3}L_{\epsilon_{3}})\Psi_{\textup{L}},\vspace{2mm}\\ \mathfrak{u}(3)&\oplus&\mathfrak{u}(1)\end{array}$ (44)
@@ IV.9 Cascade of particle symmetries: the Georgi-Glashow line
$\begin{array}[]{lll}\ell_{5}&=&R_{cd}\left(a_{c}a^{\dagger}_{d}-a_{d}a^{\dagger}_{c}\right)\vspace{2mm}\\ &+&R_{cd}^{\prime}\hskip 1.42262pti\left(a_{c}a^{\dagger}_{d}+a_{d}a^{\dagger}_{c}\right)\vspace{2mm}\\ &+&R_{c}\hskip 1.42262pti\left(a_{c}a^{\dagger}_{c}-a_{c+1}a^{\dagger}_{c+1}\right),\end{array}$ (45)
@@
$\begin{array}[]{lllll}&&\ell_{5}\Psi_{\textup{L}}=&&\vspace{2mm}\\ &&\big{(}r_{j}i\Lambda_{j}+r^{\prime}_{m}L_{\epsilon_{m}}s+irY&&\vspace{2mm}\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{1}\left(L_{\epsilon_{1}}L_{e_{4}}+L_{\epsilon_{2}}L_{e_{5}}\right)&+&it_{1}^{\prime}\left(L_{\epsilon_{1}}L_{e_{5}}-L_{\epsilon_{2}}L_{e_{4}}\right)\vspace{2mm}\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{1}^{\prime\prime}\left(L_{\epsilon_{3}}L_{e_{4}}-iL_{e_{5}}L_{e_{7}}\right)&+&it_{1}^{\prime\prime\prime}(L_{\epsilon_{3}}L_{e_{5}}+iL_{e_{4}}L_{e_{7}})\vspace{2mm}\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{2}\left(L_{\epsilon_{1}}L_{e_{1}}+L_{\epsilon_{2}}L_{e_{3}}\right)&+&it_{2}^{\prime}(L_{\epsilon_{1}}L_{e_{3}}-L_{\epsilon_{2}}L_{e_{1}})\vspace{2mm}\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{2}^{\prime\prime}\left(L_{\epsilon_{3}}L_{e_{1}}-iL_{e_{3}}L_{e_{7}}\right)&+&it_{2}^{\prime\prime\prime}(L_{\epsilon_{3}}L_{e_{3}}+iL_{e_{1}}L_{e_{7}})\vspace{2mm}\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{3}\left(L_{\epsilon_{1}}L_{e_{2}}+L_{\epsilon_{2}}L_{e_{6}}\right)&+&it_{3}^{\prime}(L_{\epsilon_{1}}L_{e_{6}}-L_{\epsilon_{2}}L_{e_{2}})\vspace{2mm}\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{3}^{\prime\prime}\left(L_{\epsilon_{3}}L_{e_{2}}-iL_{e_{6}}L_{e_{7}}\right)&+&it_{3}^{\prime\prime\prime}(L_{\epsilon_{3}}L_{e_{6}}+iL_{e_{2}}L_{e_{7}})\big{)}\Psi_{\textup{L}}\vspace{2mm}\\ &&\hskip 48.36958pt\mathfrak{su}(5)&&\end{array}$ (46)
@@
With our $\mathfrak{su}(5)$ action now established, division algebraic reflections may be employed so as to break these symmetries sequentially.
@@
$\begin{array}[]{rccll}\ell_{\textup{SM}}\Psi_{\textup{L}}=\big{(}r_{j}i\Lambda_{j}&+&r^{\prime}_{m}L_{\epsilon_{m}}s&+&irY\big{)}\Psi_{\textup{L}}.\vspace{2mm}\\ \mathfrak{su}(3)&\oplus&\mathfrak{su}(2)&\oplus&\mathfrak{u}(1)\end{array}$ (50)
@@
$\begin{array}[]{rccll}\ell_{\textup{UBZ}}\Psi_{\textup{L}}=\big{(}r_{j}i\Lambda_{j}&+&r^{\prime}_{3}L_{\epsilon_{3}}s&+&irY\big{)}\Psi_{\textup{L}}.\vspace{2mm}\\ \mathfrak{su}(3)&\oplus&\mathfrak{u}(1)&\oplus&\mathfrak{u}(1)\end{array}$ (52)
@@
$\begin{array}[]{lrcl}\ell_{\textup{UB}}\Psi_{\textup{L}}=&\big{(}r_{j}i\Lambda_{j}&+&r^{\prime}iQ\big{)}\Psi_{\textup{L}},\vspace{2mm}\\ &\mathfrak{su}(3)&\oplus&\mathfrak{u}(1)\end{array}$ (54)
@@
Readers may observe in particular that the familiar complex conjugate is what dictates the Standard Model’s final unbroken gauge symmetries. Conceivably, the Standard Model’s gauge group could have broken in many different ways, had it been paired with different Higgs sectors. Could the complex conjugate ultimately be stewarding the symmetry breaking process?
@@ V Roadmap
In this paper, we have listed five different forms of symmetry breaking steps. Namely, the multivector condition, $\Psi_{V}$, invariance under two types of octonionic reflection, ${*_{\overrightarrow{\mathbb{O}}}}$ and $*_{\overleftarrow{\mathbb{O}}}$, one quaternionic reflection, $*_{\overrightarrow{\mathbb{H}}},\vspace{.5mm}$ and one complex reflection, $*_{\overrightarrow{\mathbb{C}}}$. These four types of reflection are precisely those that can be realized in terms of the left-multiplication algebra of $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}.$
@@
Our culminating diagram results when all five symmetry breaking steps are amalgamated ($\Psi_{V}$, ${*_{\overrightarrow{\mathbb{O}}}}$, $*_{\overleftarrow{\mathbb{O}}}$, $*_{\overrightarrow{\mathbb{H}}},$ $*_{\overrightarrow{\mathbb{C}}}$). For reflections, we begin with the octonions, continue to the quaternions, and close with the complex numbers. It is interesting to note that the order $\mathbb{O}\mapsto\mathbb{H}\mapsto\mathbb{C}$ does indeed matter, otherwise one would not necessarily expect our results to hold. Please see Figure (6) at the end of this article.
Figure 6: Six well-known models of elementary particles are interconnected algebraically into a detailed particle roadmap. The “SO(10)” model (Spin(10)), the Georgi-Glashow model (SU(5)), the Pati-Salam model (PS), the Left-Right Symmetric model (LR), the Standard Model pre-Higgs-mechanism (SM), and the Standard Model post-Higgs-mechanism (UB) are each interrelated via certain algebraic constraints. The multivector condition is represented by edges labeled as $\Psi_{V}$. Division algebraic reflection constraints are represented by edges labeled as $*_{\mathbb{D}}$. Beyond these six well-known models, there are three more. For the definition of UBZ+ see equation (42); for the definition of UB+ see equation (44); for the definition of UBZ see equation (52).
From the outset, readers may notice an interesting feature of this 5-step network. Namely, certain pairs of nodes accommodate multiple pathways between them.
As a first example, we identify two compatible symmetry breaking pathways from the Spin(10) model to the Left-Right Symmetric model. One may first restrict Spin(10) to those symmetries invariant under the octonionic reflection $*_{\overrightarrow{\mathbb{O}}}$, resulting in the Pati-Salam model. Subsequently imposing invariance under ${*_{\overleftarrow{\mathbb{O}}}}$ then results in the Left-Right Symmetric model. As an alternate route, readers can bypass directly to the Left-Right Symmetric model by simply restricting Spin(10) once via $*_{\overleftarrow{\mathbb{O}}}$, as in [32].
@@
VI Outlook
With this framework of particle theories now set in place, one might wonder if there is a common thread connecting each of the symmetry breaking steps displayed here. Indeed, there is.
As pointed out in [9], the multivector condition ($\Psi_{V}$) may be seen to be equivalent to fixing a volume element in the exterior/Clifford algebraic representation of $\Psi$. Likewise, as pointed out in [3], each of the division algebraic reflection steps $\{{*_{\overrightarrow{\mathbb{O}}}},*_{\overleftarrow{\mathbb{O}}},*_{\overrightarrow{\mathbb{H}}}\}$ may be seen to be equivalent to fixing a volume element in the multiplication algebras (associated Clifford algebras) of the various division algebras. (The complex volume element is trivially preserved.) The connection of these volume elements to Hodge duality and Jordan algebras will be explored in future work.
VII Conclusion